XYZ company has the following: –Stock market price, December 31, 2018: $80 –Common stock ($5 par): $3,000,000 –Net Income for year 2018: $1,700,000 –Retained Earnings January 1, 2018: $1,900,000 If on December 31, 2018, the company wants to pay the maximum amount it can in dividends to its shareholders, what is the dividend yield ratio? Select one: a. 7.5% b. 4% c. 3.5% d. 6%
Answer #1

Dividend per share = (Net income + Retained earnings) / Outstanding common shares

= (1,700,000+1,900,000)/(3,000,000/5)

= 3,600,000/600,000

= 6 per share

Dividend yield ratio = Dividends per share/Market price per share

= 6/80

= 7.5%